Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
March 15, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Winston Harris, Michael Howell, Michael Jay, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Mark Sullivan, Evan Wack

Guest: Rachel Schipper
Review Minutes of 3/8/10
  1. Network #1 - Keith Sanders, not Keith Monroe.
  2. Software #9 - database not moved yet.
Ongoing IT Updates

Orders
  1. Order for 174 replacement computers approved for West public and out-of-warranty staff machines.
  2. An additional 200 will go out of warranty at the end of the year - we may replace them.
Relocation
  1. East campus visit last week - room layouts don't line up with windows. Rachel to check with the architect about fixing. The work room/computer storage area has a wall of windows. Will says the conference room looks good.
  2. Looks to be plenty of parking. Parking will be free for two years. Parking permit renewals come up in May. There will be a shared pass. For those going back and forth, it is recommended that you keep your campus parking pass. Additionally, the Facilities van will be available for checkout.
  3. No occupancy date - Rachel will try to obtain. Scheduled for completion on June 1st.
  4. MSL tapes - Will to contact Tara regarding disposal.
Hardware
  1. Public printing - no change.
  2. Josh is testing four computers: two with Clean Slate and two with ReturnIL. ReturnIL can be installed upon a standalone server. Clean Slate - Amy has a quote. Evan requests she send it to him. Logan also has it. On paper, Clean Slate looks a bit better. ReturnIl is a reboot.
  3. Help Desk - response to phones is positive.
  4. HIP, Site Advisor - released last week. ALA site got red on Site Advisor. Small virus incident today - two machines hit - same virus as last month. Site advisor does not work in Chrome.
  5. Video Conferencing - the company was here last week. They did not fix the microphones. Evan sent them a strong email stating our needs. Two issues: microphones and DVR. Training on the equipment scheduled for today with Tom Caswell.
  6. Altiris - waiting for the PO. Should have a positive response in the next day or so. It is making its way through the chain of the campus purchasing queue.
  7. Win7 staff image is done. Will start on the public image tomorrow. We will be installing Win7 on Judy's new Toughbook.
  8. Altiris - Ghost, Wins, our own image. Altiris innovators requested a public IP. Port at CNS - need to request that it be open. Will conduct a pre-flight conference call before they arrive.
  9. Time-limited workstations - Amy checked with Stacey last week. We are still waiting. Should be going directly to John Ingram for verification.
Software
  1. VIVO - we did make a job offer for the first position.
  2. Patron updates failed. It has 31 rather than 51,000 students. The download from the Registrar did not work.
  3. Macros - Chelsea's is almost done. Need to digitize ALEPH spreadsheet. Preservation project - on hold for Christy Shorey. Todd Chisholm has an elaborate book/magazine macro.
  4. Hours GUI is ongoing.
  5. Shruti made changes to the Staff Directory. Sent the interface to Tom Minton.
Network/Server Administration
  1. DHCP - no word from CNS - Logan will call them today. If we end up having to do it, it would be a major project. The question is if it is configured for a Pixie boot.
  2. Virtualization - Keith got the upgrade to 4.0 finished. May have SQL and two Altiris servers done by tomorrow.
  3. CNS backup - spreadsheet of costs - need to bump up. Weekly is most cost effective. It is two-thirds less than daily. Estimate that more will cover weekly backup - DLC SAN only. Monthly scan gets the biggest hits. Our biggest storage is the SAN. Moving to managed servers leaves only small servers, such as EPO. Rachel asks: can anything be cancelled to make up the difference? Tivoli cost will increase, Backup Exec license could be let go, which would replace the cost of equipment and tapes.
  4. Streaming video - pending. Cathy has a 3.8 Gig video she wants to stream. Fallback plan would be to target a 10 minute (most applicable) portion on the database. Could put it on YouTube. We have no method to stream right now. Evan will call Cathy. Evan will give this to Matthew. Rachel says there may be another way to do this - we have a donor. Working on getting the video digitized. Looking into virtualizing. Tom says another avenue may be ITunes. Could brand to the University, like a podcast. This hasn't been marketed very much.
ILS
  1. Nothing new.
  2. Training for Cataloging starts the 23rd.
  3. Logan will meet with Jason on Thursday.
Web Projects
  1. Proxy server - FCLA hosts for FAMU, not anybody else. Tom spoke with Cliff Richmond. They have their own VPN with no client. Tom spoke with him on Thursday at the Emerging Technologies meeting. Ball is in his court. Jason polled state universities using EZProxy. This will be a policy issue. Looking into a Library Proxy. VPN problem - CISCO client is flaking on VISTA and Win7. There isn't one for 64-bit. Reports of patrons that are on-campus not being able to access the database - depends on IP range provided to vendors. EZProxy is not Shibboleth. Winston talked to Jennifer about plans to Shibboleth ALEPH. It is just starting to be discussed at planning meetings. 14-digit concept will be problematic - not a uniform approach. EZProxy is sort of a Library standard. New product must use Shibboleth.
  2. Ben Walker - UF-YouBorrow - statewide borrowing to be used when the storage facility is up. There is a pilot of the project, but we are not a part of it. Should be by the end of the summer.
  3. VT Calendar - move to Logan. Should be just a copy - need the URL to change on the home page. Got a list of folks from Barbara Hood as to who will submit entries.
  4. IPhone - still testing. Basically done.
  5. CMS - 19.50 for port is still the blog server, not separate. Logan will fix.
DLC General Issues
  1. Database at CNS - Mark Sullivan cannot access.
  2. UFDC - new version release.
  3. Question about web move to CNS - IS7 wasn't done, but should be this week.
  4. Mark is working on a Google-based map search for the Aerial Photography Collection to augment GIS.
Students and Department
  1. Restrooms in East are being renovated. Will be ongoing for 2-3 weeks. Will start on the first floor. Peter & Rob (Facilities) will both be out this week. If you have any building issues, please get in touch with Brad. Will to email Brad about the fire hazard if the south tower door is closed off.